OpenStudy (lgbasallote):

I just need to know how to solve for the area of the drawing at the bottom comment

OpenStudy (anonymous):

It's significantly easier that way.

OpenStudy (lgbasallote):

so i find for the intersection right? \[2y - y^2 = y^2 - 4y\] \[2y^2 - 6y = 0\] \[2y(y - 3) = 0\] y = 0 y = 3 so these are my upper and lower limits \[\int_{0}^{3} (2y - y^2 -[y^2-4y])dy\] did i do it right?

OpenStudy (anonymous):

Yes, perfect.

OpenStudy (lgbasallote):

yay thanks :DD i think i got the curves down \m/ btw...how do you do those circles?? they're weird

OpenStudy (anonymous):

Hmm? Circles?

OpenStudy (lgbasallote):

i dont know my teacher showed examples having those buggers

OpenStudy (anonymous):

 Is that what you're askin' for?

OpenStudy (lgbasallote):

 example